# Rethinking Spectral Graph Neural Networks with Spatially Adaptive Filtering

### v3

@@ 6.5. Parameter Study
Figure 6. Sensitivity analysis for hyper-parameters: $\tau$ (blue), $\eta$ (red), $\epsilon$ (green), and $L$ (orange) from top to bottom rows.
7. Conclusion and Future Work
This paper introduces a fresh spatial perspective on spectral GNNs, shedding light on their interpretability. We reveal that spectral GNNs fundamentally leads the original graph to an adapted new one, which exhibits non-locality and accommodates signed edge weights. This insight leads to our proposed Spatially Adaptive Filtering (SAF) framework, enhancing spectral GNNs for more effective and versatile graph representation learning. While this study focuses on a node-level investigation, it raises intriguing questions about the implications of spectral GNNs at the graph-level in the spatial domain. Future work could expand this examination by exploring the interplay between spatial and spectral domains from a more comprehensive graph-level viewpoint.
Appendix A Proofs
A.1. Proof of Proposition 2
Proof.
A.2. Proof of Proposition 3
Proof.
A.3. Proof of Proposition 1
Proof.
$p(x)$ $=\sum_{k=0}^{K}\psi_{k}\binom{K}{k}(1-x)^{K-k}x^{k}\leq\psi_{\text{max}}\sum_{k=0}^{K}\binom{K}{k}(1-x)^{K-k}x^{k}$
$=\psi_{\text{max}}(1-x+x)^{K}=\psi_{\text{max}}.$
Therefore, we have $p(x)\leq\max\{\psi_{k}\}_{k=0}^{K}$ for all $x\in[0,1]$. ∎
Table 3. Statistics of real-world datasets. $F$ and $C$ denotes the number of features and classes. $\Delta$ represents graph diameter referring to the longest geodesic distance between nodes on the graph. For Penn94, due to multiple subgraphs, we report $\Delta$ of the largest connected component.
[…]
Chameleon 2,277 31,371 2,325 5 11 0.23 0.04 0.03
Squirrel 5,201 198,353 2,089 5 10 0.22 0.03 0.01
Texas 183 279 1,703 5 8 0.11 0.00 -0.23
Cornell 183 277 1,703 5 8 0.30 0.02 -0.08
Actor 7,600 26,659 932 5 12 0.22 0.01 0.00
Cora 2,708 5,278 1,433 7 19 0.81 0.77 0.77
Citeseer 3,327 4,552 3,703 6 28 0.74 0.63 0.67
Pubmed 19,717 44,324 500 5 18 0.80 0.66 0.69
Minesweeper 10,000 39,402 7 2 99 0.68 0.01 0.01
Tolokers 11,758 519,000 10 2 11 0.59 0.18 0.09
Amazon-ratings 24,492 93,050 300 5 46 0.38 0.13 0.14
Roman-empire 22,662 32,927 300 18 6,824 0.05 0.02 -0.05
Penn94 41,554 1,362,229 5 2 8 0.47 0.05 0.02
Appendix B Eigendecomposition
Eigendecomposition breaks down a matrix into its eigenvalues and eigenvectors, offering insights into matrix properties, especially for the graph Laplacian. Despite computational demands, this technique has attracted surging interest in the graph learning community due to its theoretical richness, and it can be practically expedited for larger graphs using Lanczos (Lanczos, 1950) and Sparse Generalized Eigenvalue (Cai et al., 2021) algorithms. Recent advancements also underscore its value in various applications such as graph positional encoding (Belkin and Niyogi, 2003; Dwivedi et al., 2020), spectral graph convolution (Liao et al., 2019), graph domain adaptation (You et al., 2022), and graph robustness (Chang et al., 2021). For example, Laplacian eigenvectors have been widely used in identifying global position of nodes in the graph (Wang et al., 2022), particularly in recent popular graph transformers (Kreuzer et al., 2021; Kim et al., 2022; Rampášek et al., 2022), enhancing their expressiveness. Innovations like SignNet, BasisNet (Lim et al., 2022), and Sign Equivariant (Lim et al., 2023) have further optimized the processing of these eigenvectors. When exploring the expressive power of GNN models, Specformer (Bo et al., 2023) employs eigendecomposition for learning set-to-set graph filters, TEDGCN (Yan et al., 2023) leverages it for adaptive weighting of eigengraphs, and FE-GNN (Sun et al., 2023) taps into singular value decomposition (SVD) for graph feature expansion. In line with these developments, our method, SAF, also utilizes eigendecomposition to explicitly create a new graph, enabling efficient non-local aggregation with signed weights to tackle long-range dependency and graph heterophily.
Appendix C Dataset Information
We conduct experiments on 13 real-world datasets from various domains. The detailed statistics are summarized in Table 3. Alongside standard data attributes, we also provide the longest geodesic (shortest-path) distance between graph nodes for better illustrating the non-local property that we investigate in Figures 1 and 7. Moreover, we adopt three metrics - edge homophily (Zhu et al., 2020b) $\mathcal{H}$, class homophily (Lim et al., 2021) $\mathcal{H}_{\text{class}}$, and adjusted homophily (Platonov et al., 2023a) $\mathcal{H}_{\text{adjusted}}$ to assess the graph’s homophily ratio, which ranges from 0 (high heterophily) to 1 (high homophily). While the first is a commonly used index, the latter two, considering class variability and potential imbalance, have been recently introduced for more accurate estimation. For our main text analysis regarding the adapted new graph, we primarily rely on the edge homophily metric, defined as =|$\mathcal{H}=|\{(v_{i},v_{j})|(v_{i},v_{j})\in\mathcal{E}\wedge\mathbf{y}_{i}=\mathbf{y}_{j}\}|/|\mathcal{E}|$, given its simplicity and wide usage. In certain compact sections of this paper, we use four-letter abbreviations for dataset names.
Appendix D Experimental Details
In this section, we provide experimental details for reproducibility. As He et al. (2022) have made a comprehensive evaluation and share the same experimental protocol with us, we directly leverage their results for models: MLP, GCN, APPNP, ARMA, GPR-GNN, BernNet, ChebNetII, GCNII and PDE-GCN on datasets including Cham., Squi., Texas, Corn., Actor, Cora, Cite., and Pubm.. For JacobiConv, LON-GNN, and OptBasisGNN, we also report their results from corresponding papers (Wang and Zhang, 2022; Tao et al., 2023; Guo and Wei, 2023b). Other experiments are performed on a machine equipped with an NVIDIA GeForce RTX 3090 (24GB) and an Intel(R) Xeon(R) Gold 5218R CPU @ 2.10GHz (20 cores).
D.1. Baseline Implementations
In our experiments, we leverage the Pytorch Geometric library (Fey and Lenssen, 2019) implementations for GCN and APPNP. For MLP, we include a sequence of linear layers, each of which is followed by batch normalization, ReLU activation, and dropout. The number of MLP layers are tuned from 1 to 5. For the remaining baselines, we resort to their publicably released code. Besides, for the work (Li et al., 2022), only the most effective model variant, GloGNN++, is included in our experiments. Models like Geom-GCN (Pei et al., 2020) and Non-Local GNNs (Liu et al., 2021), surpassed by these SOTA methods, are excluded from our comparison.
D.2. Hyper-parameters Setting
[…]
Figure 7. Distributions of original hop distance between adjacent nodes in the adapted new graph on additional seven datasets.
Appendix E More Experiments
E.1. Empirical Evidence of the New Graph
Additional empirical support for the newfound graph’s properties is presented in Figures 7 and 8.
Figure 8. Additional visualizations for illustrating the newfound graph’s ability in discerning label consistency.
E.2. Attention Trends
For the analysis of attention trends, we provide visualizations on more datasets in Figure 9. It is noteworthy that, despite Pubmed being a homophilic graph with $\mathcal{H}=0.80$, non-local aggregation maintains a pivotal role in our SAF, diverging from the patterns observed on the Cora dataset as shown in Figure 5(b). This discrepancy ownes much to the sizable node count within Pubmed, where the capability of non-local aggregation in capturing long-range dependencies proves advantageous for improving model performance.
(a) Chameleon
(b) Pubmed
Figure 9. Additional changing trends of the averaged attention values with respect to training epochs.
E.3. New Benchmarks for Graph Heterophily
For a more extensive evaluation across various domains, we also test SAF on five recently introduced datasets, including Mine., Tolo., Amaz., Roma., and Penn94 (Platonov et al., 2023b; Lim et al., 2021). In this context, we draw comparisons solely with MLP, GCN, APPNP, along with six GNN models that have previously shown promising results in prior tasks, namely GPR-GNN, BernNet, ChebNetII, JacobiConv, FE-GNN, and ClenshawGCN. Table 4 lists the average classification accuracies, obtained over random splits provided by (Platonov et al., 2023b; Lim et al., 2021), with a distribution of 50%/25%/25% for training/validation/testing. In summary, SAF achieves significant performance gains of 12.84% and 4.10% on Mine. and Tolo. datasets, respectively, while maintaining competitiveness on the other three.
Table 4. Evaluations on new heterophilic graph datasets
@@ E.3. New Benchmarks for Graph Heterophily
Table 5. Full-supervised node classification accuracy (%). SAF-Cheb refers to SAF implementation using Chebyshev polynomials.
@@
E.4. SAF with ChebNetII as Base Model
To expand the versatility of our SAF framework, we introduced ChebNetII (He et al., 2022) as an alternative base model, chosen for its adherence to the non-negative constraint, critical in our model design as stated in Section 5.1. The rationale behind this choice is ChebNetII’s use of Chebyshev interpolation for learning Chebyshev polynomials, where the constraint can be ensured by keeping its learnable parameters $\{\gamma_{j}\}_{j=0}^{K}$ non-negative. Our experiments, as shown in Table 5, confirm that SAF can significantly enhances ChebNetII’s performance, underscoring the framework’s flexibility with different spectral filters. Interestingly, we observed that SAF, utilizing Bernstein polynomials (SAF-Bern), slightly surpasses its performance with Chebyshev polynomials (SAF-Cheb) in most datasets. The margin of improvement over the base model is also more pronounced with SAF-Bern. This phenomenon could be attributed to the $g_{\phi}(\lambda)\leq 1$ constraint within SAF (refer to Section 5.1), necessitating the rescaling of filter functions by their maximum values. For Bernstein polynomials, this maximum is readily obtained as the largest polynomial coefficient $\max\{\phi_{k}\}_{k=0}^{K}$, as per Proposition 3. However, for Chebyshev polynomials, the best theoretical upper bound is the sum of absolute coefficients, $\sum_{k=0}^{K}|\phi_{k}|$, which is comparatively less precise. This difference may impact the quality of graph construction and, subsequently, the model’s performance. Exploring these nuances will be a focal point of our future research.
Table 6. Time overheads (s) / Space overheads (MB). For large-scale graph Penn94 (with 41,554 nodes and 1,362,229 edges), we only employ a partial eigendecomposition with 100 extremal eigenvalues to balance between model effectiveness and efficiency.
@@ E.4. SAF with ChebNetII as Base Model
BernNet 8.36/72 13.74/232 3.92/5 4.16/5 4.88/292 5.24/64 5.52/152 6.06/1546 24.05/1902
ChebNetII 22.82/72 30.73/231 11.47/5 9.64/5 14.88/291 19.96/63 16.14/152 36.91/1584 41.67/1850
ClenshawGCN 15.58/98 26.21/398 2.98/5 4.13/5 4.43/313 3.72/68 12.48/152 24.97/1833 191.30/3017
SAF 11.55/112 18.78/440 4.38/5 4.70/5 5.36/733 6.04/120 6.12/237 18.43/4515 23.49/8491
Decomposition 0.58/141 1.59/540 0.02/1 0.02/1 3.93 /1206 1.00/140 0.77 /239 21.34/7641 4.76/4
Table 7. Average running time per epoch (ms) / average total running time (s).
@@
Method Cham. Squi. Texas Corn. Actor Cora Cite. Pubm.
BernNet 14.60/8.36 17.40/13.74 16.10/3.92 14.30/4.16 14.40/4.88 14.90/5.24 16.20/5.52 15.10/6.06
ChebNetII 39.30/22.82 40.70/30.73 42.30/11.47 39.60/9.64 41.80/14.88 39.20/19.96 40.70/16.14 40.60/36.91
JacobiConv 11.10/6.54 11.20/10.01 11.60/3.86 11.20/2.93 11.70/4.22 11.50/5.54 11.30/6.02 11.40/8.47
NodeFormer 135.00/58.96 135.10/79.66 49.10/14.29 67.60/18.89 158.20/66.20 56.70/19.25 73.10/32.00 150.60/68.57
GloGNN++ 53.60/35.63 123.70/68.31 14.80/4.47 9.80/3.00 204.80/73.13 89.60/32.68 49.60/12.35 6369.20/5266.53
ClenshawGCN 20.6/15.58 39.7/26.21 11.0/2.98 14.5/4.13 13.7/4.43 11.0/3.72 29.7/12.48 68.1/24.97
SAF-$\epsilon$ 30.70/19.11 59.50/54.71 27.30/6.95 29.70/7.86 121.20/33.14 30.20/10.32 32.00/17.19 2054.80/837.10
SAF 19.30/11.55 20.90/18.78 17.70/4.38 18.60/4.70 18.80/5.36 18.90/6.04 18.10/6.12 43.30/18.43
Table 8. Ablation study of SAF framework.
Our Variant Cham. Squi. Texas Corn. Actor Cora Cite. Pubm. Mine. Tolo. Amaz. Roma.
SAF-$\epsilon$ 74.84 64.00 94.75 93.28 42.98 89.87 81.45 91.52 90.59 79.45 50.36 74.89
SAF 75.30 63.63 94.10 92.95 42.93 89.80 80.61 91.49 90.54 79.38 50.49 74.87
SAF w/o Atte. 75.01 62.62 89.18 86.07 41.53 88.80 80.37 91.24 89.45 76.18 49.98 74.25
SAF w/o Spec. 73.55 55.70 90.49 88.20 41.06 88.03 78.87 90.12 89.45 78.23 49.13 71.85
SAF w/o Spat. 68.53 51.39 92.62 92.13 41.71 88.51 80.08 88.51 77.75 75.35 49.84 74.56
E.5. Time and Space Overheads
@@
Eigendecomposition. Our SAF framework pre-computes eigendecomposition once per graph and reuses it in Eq. (6). This aspect is crucial, as the forward-pass cost in model training often exceeds the preprocessing expense of eigendecomposition. To empirically validate this, we compare the time overheads of eigendecomposition with the training times of various models in Table 6. It is evident that for most datasets, the time consumed by decomposition is significantly less than the time required for model training. For medium-sized graphs such as Pubmed, while the full decomposition time exceeds that of BernNet, it still maintains efficiency against more advanced GNNs (ChebNetII, ClenshawGCN, SAF). Moving to the large-scale graph, Penn94, where only partial eigendecomposition with 100 extremal eigenvalues is considered, the computation time is markedly reduced compared to all models mentioned.
@@
Model Comparison. To evaluate the efficiency of our model, we compare the running times of our proposed variants, SAF and SAF-$\epsilon$, against four notable spectral GNNs (BernNet, ChebNetII, JacobiConv), two non-local GNNs (NodeFormer, GloGNN++), and one unified GNN model (ClenshawGCN), as detailed in Table 7. One can observe that SAF, while slightly slower than its base model, BernNet, due to the integration of non-local spatial aggregation, remains more efficient than or comparable to other SOTA methods. On the other hand, SAF-$\epsilon$ costs more time (but still faster than the non-local GNN, GloGNN++), a consequence of its quadratic complexity from non-local sparsification. However, this trade-off allows SAF-$\epsilon$ to produce an high-quality new graph, better capturing long-range interconnections among nodes and addressing graph heterophily.

### v5

@@ VI-D Analysis of Attention Trends
Figure 6: Sensitivity analysis of hyper-parameters: $\tau$, $\eta$, $\epsilon$, and $L$ from top to bottom rows.
VI-E Ablation Study
This section aims to validate our designs by comparing SAF with its three ablated variants – SAF w/o Atte., SAF w/o Spec., and SAF w/o Spat. – in full-supervised node classification. Specifically, Atte., Spec., and Spat. respectively refers to: attention mechanism in “Node-wise Prediction Amalgamation”, “Non-negative Spectral Filtering”, and “Non-local Spatial Aggregation”. For SAF w/o Atte., we remove the attention mechanism and equally blend predictions from different domains. SAF w/o Spec. abandons the spectral filtering phase, practically setting $\boldsymbol{\kappa_{f}}=\mathbf{0},\boldsymbol{\kappa_{a}}=\mathbf{1}$. As the SAF w/o Spat. configuration is equivalent to BernNet model, the corresponding results are posted directly. From Fig. 5, we can draw several insights: 1) The impact of Atte. module on our SAF varies by datasets, e.g., on Chameleon and Squirrel, showing a slight performance reduction upon its removal. This observation aligns with our observation that their optimal attention values are close to an even split, as suggested in Figures 4(a). Conversely, Cora dataset exhibits a notable drop, due to its optimal attention weights being far from even, as depicted in Fig. 4(b); 2) Spectral filtering (Spec. module) remains vital for discriminative node representation learning. Specifically, the quality of the adapted new graph fundamentally hinges on the graph spectral filters’ training, as underscored by their theoretical interaction in Eq. (6). Practically, the absence of spectral filtering markedly reduces model accuracy, confirming its importance in SAF; 3) This visualization not only reaffirms the pivotal role of the non-local aggregation (Spat. module), but also underscores its position as the most crucial component in advancing spectral GNNs within the SAF framework.
TABLE IV: Evaluations on new heterophilic graph datasets.
@@ VI-E Ablation Study
VI-F Parameter Study
This section presents the sensitivity analysis of hyper-parameters including $\tau$, $\eta$, $\epsilon$, and $L$. Fig. 6 visualizes how varying these parameters within a broad range influences learning performance, showcasing our model’s robust stability over diverse settings. Beyond empirical observation, we also provides deeper insights into parameter understanding and rationalizes the chosen ranges for parameter searching: 1) The scaling parameter $\tau=\frac{\alpha}{1-\alpha}$, crucial in new graph construction in Eq. (6), stems from the trade-off parameter $\alpha\in(0,1)$ within the graph optimization problem in Eq. (3). While theoretically we have $0=\frac{0}{1-0}<\tau<\frac{1}{1-1}=\infty$, practical considerations for extracting structural information suggest a larger penalty on the trace objective term $\mathrm{tr}(\mathbf{Z}^{T}\gamma_{\theta}(\hat{\mathbf{L}})\mathbf{Z})$, i.e., keeping $\alpha<0.5$, thereby limiting $\tau<\frac{0.5}{1-0.5}=1$. This rationale substantiates our selection of $\tau$ within the set $\{0.1,0.2,...,1\}$ as stated in Section VI-A4, aligning with the observed optimal performance in Figures 6(a)-(e). When addressing graphs with noisy structure, we may adjust the upper limit of $\alpha$ to $t\in(0,1)$, setting $\tau$’s maximum possible value to $\frac{t}{1-t}$. For graph benchmarking evaluations in this work, where extracting structural information is important, we practically set $t=0.5$; 2) For the non-local aggregation layer number $L$, a noticeable decline in model performance is observed when $L$ exceeds 10. This is attributed to the non-local nature of our new graph, which facilitates efficient information exchange between nodes. Exceeding a certain number of layers may potentially lead to oversmoothing, where there is an overemphasis on global information, thus degrading model performance. However, choosing the number of layers within a reasonable range generally ensures consistent and impressive model performance, as verified in Figures 6(j)-(l).
TABLE V: Full-supervised node classification accuracy (%) while implementing SAF upon ChebNetII.
@@ VI-F Parameter Study
VI-G New Benchmarks for Graph Heterophily
For a more extensive evaluation across various domains, we also test SAF on five recently introduced datasets, including Mine., Tolo., Amaz., Roma., and Penn94 [84, 67]. In this context, we draw comparisons solely with MLP, GCN, APPNP, along with seven GNN models that have previously shown promising results in prior tasks, namely GPR-GNN, BernNet, ChebNetII, JacobiConv, NodeFormer, GloGNN++ and FE-GNN. Table IV lists the average classification accuracies, obtained over random splits provided by [84, 67], with a distribution of 50%/25%/25% for training/validation/testing. In summary, SAF achieves significant performance gains of 12.79% and 4.03% on Minesweeper and Tolokers datasets, respectively, while maintaining competitiveness on the others.
VI-H SAF with ChebNetII as Base Model
To expand the versatility of our SAF framework, we introduced ChebNetII [15] as an alternative base model, chosen for its adherence to the non-negative constraint, critical in our model design as stated in Section V. The rationale behind this choice is ChebNetII’s use of Chebyshev interpolation for learning Chebyshev polynomials, where the constraint can be ensured by keeping its learnable parameters $\{\gamma_{j}\}_{j=0}^{K}$ non-negative. Our experiments, as shown in Table V, confirm that SAF can significantly enhances ChebNetII’s performance, underscoring the framework’s flexibility with different spectral filters. Interestingly, we observed that SAF, utilizing Bernstein polynomials (SAF-Bern), slightly surpasses its performance with Chebyshev polynomials (SAF-Cheb) in most datasets. The margin of improvement over the base model is also more pronounced with SAF-Bern. This phenomenon could be attributed to the $g_{\phi}(\lambda)\leq 1$ constraint within SAF (refer to Section V), necessitating the rescaling of filter functions by their maximum values. For Bernstein polynomials, this maximum is readily obtained as the largest polynomial coefficient $\max\{\phi_{k}\}_{k=0}^{K}$, as per Proposition 3. However, for Chebyshev polynomials, the best theoretical upper bound is the sum of absolute coefficients, $\sum_{k=0}^{K}|\phi_{k}|$, which is comparatively less precise. This difference may impact the quality of graph construction and, subsequently, compromise the model’s performance. Exploring these nuances will be a focal point of our future research.
Figure 7: Improved classification accuracy by GloGNN++, LRGNN and our SAF (from left to right) over the vanilla GCN under semi-supervised setting (2.5%/2.5%/95%).
VI-I Supervision Dependence in GloGNN++ and LRGNN
This subsection empirically verifies the heavy reliance on label supervision for both GloGNN++ [56] and LRGNN [86] models. In Fig. 7, we compare the improved classification accuracy of these models and our SAF upon the vanilla GCN under a semi-supervised setting. The figure illustrates that, despite competitive results under dense supervision (Table III), the performance of both GloGNN++ and LRGNN deteriorates, demonstrating negative optimization on many datasets, when the label rate drops approximately from 60% to 2.5%. This degradation highlights their dependency on high label rates for effective learning. In contrast, our SAF constantly delivers good performance regardless of the supervision level, showing its effectiveness in both high and low label rate scenarios.
TABLE VI: Time overheads (s).
@@ VI-I Supervision Dependence in GloGNN++ and LRGNN
# Nodes 2,277 5,201 183 183 7,600 2,708 3,327 19,717 41,554
# Edges 36,101 217,073 309 295 33,544 5,429 4,732 44,338 1,362,229
BernNet 8.36 13.74 3.92 4.16 4.88 5.24 5.52 6.06 24.05
ChebNetII 22.82 30.73 11.47 9.64 14.88 19.96 16.14 36.91 41.67
FE-GNN 3.99 45.89 1.03 0.84 0.98 2.54 2.06 6.78 7.33
NodeFormer 58.96 79.66 14.29 18.89 66.20 19.25 32.00 68.57 122.91
GloGNN++ 35.63 68.31 4.47 3.00 73.13 32.68 12.35 5266.53 3614.37
SAF 11.55 18.78 4.38 4.70 5.36 6.04 6.12 18.43 23.49
Decomposition 0.58 1.59 0.02 0.02 3.93 1.00 0.77 21.34 4.76
TABLE VII: Space overheads (MB).
@@
Method Cham. Squi. Texas Corn. Actor Cora Cite. Pubm. Penn94
BernNet 72 232 5 5 292 64 152 1546 1902
ChebNetII 72 231 5 5 291 63 152 1584 1850
FE-GNN 1337 6919 23 10 416 302 740 2213 5854
NodeFormer 1522 3965 15 37 775 480 764 2119 3056
GloGNN++ 290 1525 5 5 2471 331 607 17892 25260
SAF 112 440 5 5 733 120 237 4515 8491
Decomposition 141 540 1 1 1206 140 239 7641 4
VI-J Time and Space Overheads
VI-J1 Eigendecomposition
@@
Our SAF framework pre-computes eigendecomposition once per graph and reuses it in Eq. (6) in both training and inference. This aspect is crucial, as the forward-pass cost in model training often exceeds the preprocessing expense of eigendecomposition. To empirically validate this, we compare the time overheads of eigendecomposition with the training times of various models in Table VI. Overall, we have following observations: 1) For datasets with a small number of nodes, the time consumed by decomposition is significantly less than the time required for model training; 2) For medium-sized graphs such as Pubmed, while the full decomposition time exceeds that of BernNet, it still maintains efficiency against more advanced GNNs such as ChebNetII, NodeFormer and GloGNN++; 3) Moving to the large-scale graph, Penn94 (with 41,554 nodes and 1,362,229 edges), where only partial eigendecomposition with 100 extremal eigenvalues is considered, the computation time is markedly reduced compared to all the models. For space overheads in Table VII, similar patterns can be observed.
VI-J2 Model Comparison
@@
In Table VI, we compare the running times of our SAF against two notable spectral GNNs (BernNet, ChebNetII), two non-local GNNs (NodeFormer, GloGNN++), and one unified GNN model (FE-GNN). Generally, one can observe that SAF, while slightly slower than its base model, BernNet, due to the integration of non-local spatial aggregation, remains more efficient than or comparable to other SOTA methods, particularly those also employing non-local approaches such as NodeFormer and GloGNN++.
VII Conclusion
This study presents a cross-domain analysis on GNN models, offering a fresh perspective by rethinking spectral GNNs from a spatial lens. We reveal that spectral GNNs fundamentally leads the original graph to an adapted new one, which exhibits non-locality and accommodates signed edge weights to reflect label consistency among nodes. This insight leads to our proposed Spatially Adaptive Filtering (SAF) framework, enhancing spectral GNNs for more effective and versatile graph representation learning. While SAF adeptly captures long-range dependencies and addresses graph heterophily, we acknowledge two limitations of this work: 1) the non-negative constraints of the proposed SAF on graph filters might limit model expressiveness, indicating room for theoretical refinement; 2) although this study focuses on a node-level investigation, it raises intriguing questions about the implications of spectral GNNs at the graph-level in the spatial domain. Future work could expand this examination by relaxing theoretical constraints or exploring the cross-domain interplay from a broader graph-level viewpoint.
Acknowledgments
The work was supported by the following: National Natural Science Foundation of China under No. 92370119, and 62376113.